### v2

# Sharp Tools: How Developers Wield Agentic AI in Real Software Engineering Tasks

Abstract
Software Engineering Agents (SWE agents) can autonomously perform development tasks on benchmarks like SWE Bench, but still face challenges when tackling complex and ambiguous real-world tasks. Consequently, SWE agents are often designed to allow interactivity with developers, enabling collaborative problem-solving. To understand how developers collaborate with SWE agents and the communication challenges that arise in such interactions, we observed 19 developers using an in-IDE agent to resolve 33 open issues in repositories to which they had previously contributed. Participants successfully resolved about half of these issues, with participants solving issues incrementally having greater success than those using a one-shot approach. Participants who actively collaborated with the agent and iterated on its outputs were also more successful, though they faced challenges in trusting the agent’s responses and collaborating on debugging and testing. These results have implications for successful developer-agent collaborations, and for the design of more effective SWE agents.
Index Terms:
Agent-based systems, software development tools, collaboration, communication, developer experience
I Introduction
Software developers are increasingly incorporating generative artificial intelligence (AI) tools into their work to boost productivity[1]. Initially, tools like inline code completions and chat assistants have aided developers by generating code snippets and providing on-demand answers. More recently, there has been a surge in the adoption of software engineering agents (SWE agents)—tools specifically designed to autonomously perform complex development tasks with the capability to iteratively refine their outputs based on environmental feedback. They not only generate code, but can also execute tool calls and act on these tools’ results [2].
SWE agents have shown promising results on established benchmarks of software development tasks, such as fully autonomously solving many of the GitHub issues in SWE-Bench [3]. However, they still struggle to solve some issues due to several factors, such as the complexity and ambiguity of the issues, lack of tacit knowledge, and limited access to the full development environment [4, 5]. To overcome these limitations, many SWE agents are designed to interact with developers, enabling a collaborative approach to problem-solving. This interaction allows developers and agents to share responsibilities such as understanding the codebase, localising, coding, debugging, testing, and reviewing each other’s work.
Figure 1 shows an example of one such SWE agent – Cursor Agent in the Cursor integrated development environment (IDE) – responding to a developer’s prompt that has been copied from a GitHub bug report [6]. In the top part of the figure, the agent searches the codebase, describes the issue based on this search, and tries to implements a fix. After prompting, the agent then adds a unit test and runs it.
While reviewing the agent’s changes, the developer notices a functional call erroneously changed by the agent (Figure 2, bottom), and rejects that part of the change, before reviewing the remaining 5 changes.
There is reason to believe that collaborations between agents and developers will not be smooth and painless. For non-agentic tools, prior research has found that developers sometimes struggle with understanding and trusting the AI’s output, hindering effective collaboration [7, 8, 9]. Is this also the case for SWE agents? On one hand, unlike non-agentic tools that generate a single response and act as a black box, agents can provide developers with information about their process by showing their work as they go. Further, their capability to make decisions autonomously allows for developers to express higher-level goals rather than low-level instructions. On the other hand, the broader scope of agents’ decisions and subsequent actions can exacerbate trust issues and place a higher cognitive load on the developer of understanding both the agent’s reasoning trajectory as well as mutations it performs to the codebase. The answer is yet unclear, due to a lack of research on real-world usage of SWE agents [2].
Figure 1: Example usage of the Cursor Agent
Figure 2: A suggested code edit diff by the agent.
In this paper, we begin to address this gap by observing 19 open-source software developers use SWE agents to solve real-world open issues of their choosing. Through this study, we aim to address the following research questions:
  1. RQ1:
    How do software developers collaborate with SWE agents to resolve open issues in active codebases?
  2. RQ2:
    What are the communication barriers to collaborations between software developers and SWE agents when addressing open issues in active codebases?
  3. RQ3:
    What factors influence participants’ success when collaborating with a SWE agent to solve open issues in active codebases?
In answering these questions, this study contributes the first empirical investigation of developer-agent collaboration in resolving real-world issues.
II Related Work
II-A Agentic Tools for Software Development
A single definition of an agentic tool has not yet been clearly established and evolves as new tools are introduced [10]. Yang and colleagues state that a language model ‘acts as an agent when it interacts with an environment by iteratively taking actions and receiving feedback’ [11], while Jin and colleagues mention that the ability to ‘select the optimal solution from multiple homogeneous results’ is fundamental to an agentic tool [12]. Aside from generating textual responses, agents can perform actions such as searching for files, modifying code, and running terminal commands by calling tools. Further, agents can iterate on their own output, enabling them to complete tasks without the need of human intervention [2].
In software development, there has been a rapid rise in the number of agentic tools in research and in practice [10]. Many of these tools aim to solve end-to-end software engineering tasks, as reflected in the evolution of benchmarks, such as SWE-bench [3], a set of real previously-solved GitHub issues, and LiveCodeBench [13], which adds new issues over time. Similarly, SWE-Lancer [4] uses real-world freelance software engineering tasks to evaluate AI tools. While these benchmarks aim to evaluate real-world performance, they all test the autonomous performance of agentic tools, rather than their performance with human-in-the-loop. As Liu and colleagues report, most research-based agentic systems target maximum automation, where human involvement is limited to specification of the problem statement [2].
At the same time, several agentic tools have been integrated into IDEs, such as VSCode Agent Mode [14], Windsurf Cascade [15] and Cursor Agent [16]. Unlike fully autonomous agentic systems, these tools are designed to operate with a human-in-the-loop. Although such tools are rapidly emerging and evolving, at the time of this writing, we know of no empirical research about how developers use such in-IDE agentic tools. This paper aims to fill this gap.
II-B Developer Experiences with AI Tools
Prior research has also explored the strategies that developers apply when using AI [17], especially prompt engineering strategies for developers [18, 19, 20]. Barke and colleagues found that developers used GitHub Copilot code completions and chat in different ways at different stages of solving an issue [8]. Prior work has also found that several development tasks are time-consuming or challenging when working with AI developer tools, including communicating the required context to AI tools [8, 17], verifying AI-generated code recommendations [21], and debugging incorrect AI-generated code [9]. Developer studies are mixed about whether AI increases developer efficiency [9] or decreases it [22]. Studies have also found that developers often use AI tools as learning aids and information sources [9, 23], and have concerns about the quality and reliability of generated code [21, 7], including code security  [9, 22]. While these studies have richly explored developer interactions with AI code completion [21, 9, 8, 24] and AI-based chat [25, 26, 27, 28], there is a dearth of developer experience studies on the use of SWE agents. Yet, SWE agents offer a unique interaction paradigm, differing from non-agentic tools in their capability to perform complex tasks autonomously and take feedback from the environment [2].
Such capabilities are also found in software bots, which can make asynchronous edits to codebases. Prior developer experience research has found that these tools sometimes perform unsolicited actions and produce too much information [29, 30]. Largely pre-dating the current wave of LLM-based AI, such bots typically operate on strict instructions; in contrast, SWE agents can make decisions autonomously, potentially complicating effective collaboration [31].
Prior work has argued for the importance of building design guidelines for agents to facilitate effective interaction mechanisms [32, 2]. Recent work has begun to address this gap for SWE agents. While Epperson and colleagues discussed the challenges developers face when debugging autonomous multi-agent systems [33], Pu and colleagues studied the effect of proactivity in developers’ interactions with agentic tools, reporting that contextual awareness and salience of the agent’s actions are critical to avoid workflow disruptions [34]. We build on this line of work by studying interactions with SWE agents in real-world tasks.
III Study Design
To answer our research questions, we aimed to observe developers collaborating with a SWE agent in a highly ecologically valid setting. To achieve this, we asked 19 professional software developers to use the Cursor Agent to resolve issues in a codebase to which they have previously contributed.
III-A Tool Selection
In selecting the appropriate SWE agent for our study, we considered the spectrum of human-AI collaboration capabilities, ranging from highly autonomous agents to those that enable robust interaction with developers. We reviewed several agents, starting with fully autonomous options like AutoCodeRover [35], and moving to agents that allow a moderate level of human-AI collaboration. These moderate collaboration tools often operate outside the IDE and primarily rely on user interaction through prompting, such as Devin [36], Aider [37],OpenHands [38] and GitHub Copilot Workspaces [39]. Finally, we considered agents that operate within an IDE, facilitating high collaboration not only through prompting but also by providing developers access to the full development environment. Examples include Cursor Agent [16], VSCode Agent Mode [14], Windsurf Cascade [15], Cline [40], and Amazon Q Developer [41].
Our focus was on IDE-based agents, as they offer more flexibility for developers to choose whether to rely fully on the agent or closely collaborate on tasks. Moreover, IDE-based AI tools like GitHub Copilot, Cursor, and Windsurf have gained significant popularity in recent years.
In assessing these tools’ capabilities, we found that Amazon Q was lacking features present in the other tools, such as output explanations, where an agent provides a textual description of the changes it made and the logic behind them. VSCode Agent Mode was newly released at the time of our study and not yet as mature as other options. Although Cline is a capable SWE agent, the grey literature suggests that it has a ‘steeper learning curve, less polished UI than Cursor or Windsurf’, [^1] https://medium.com/@pahwar/cline-vs-windsurf-vs-pearai-vs-cursor-2025s-top-ai-coding-assistants-compared-2b04b985df51 making it less appropriate for a study with limited time for participants to learn the tool.
This left Cursor Agent and Windsurf Cascade as our final candidates. We conducted pilot testing using both tools with a small convenience sample of developers and found the tools comparable in capabilities and usability. However, Cursor Agent offered one feature that made it particularly suitable for our study: automatic incorporation of users’ currently open files and cursor position as context for the agent. Therefore, we selected Cursor Agent as our study tool. Participants used Cursor v 0.47 with Claude 3.5 Sonnet.
III-B Tasks
We sought software engineering tasks for this study that:
  • reflected developers’ typical day-to-day tasks;
  • participants were motivated to solve thoroughly;
  • could be completed in less than an hour, enabling us to analyze end-to-end agent-participant collaboration; and
  • would not use proprietary code, which participants may not be allowed to show us as researchers or to send to a third-party AI tool.
Thus, we asked participants to work on recent open issues (such as bugs or feature requests) in open-source repositories that they had worked on before. We applied the following guidelines to select suitable repositories that were:
  • of significant size, that is, with more than 50 source code (i.e., non-documentation and non-configuration) files;
  • actively maintained (i.e., had commits in the prior month);
  • software centric, thus excluding repositories such as a collection of study materials or tutorials [42]; and
  • starred over 500 times on GitHub, an indicator of the maturity of the repository [43].
Additionally, we selected repositories that were part of two major GitHub organizations that are managed by our company. This enabled us to select participants based on internal-only information about repository contributors, like seniority and geographic location, and to communicate more easily with those contributors.
While most participants selected their first task to complete based on the list of open issues in their project, some participants selected tasks to work on ahead of time. We further asked participants to choose issues that could be solved based on code edits, rather than configuration or documentation fixes, to ensure the agent could provide meaningful assistance.
If participants had time to work on multiple tasks during the session, after they completed their first task, we collaborated with them to choose subsequent issues to ensure diversity in:
  • task type (e.g., if the first issue was a bug, we encouraged participants to next select a feature request), and
  • difficulty and scope (e.g., if the first issue was completed easily, we encouraged participants to next select an issue that seemed more challenging).
TABLE I: Participant Demographics
Dimension Details
Professional Role Software Engineer I/II: 6; Senior Software Engineer: 7; Principal Software Engineer/Manager: 5; Associate Consultant: 1
Professional Experience 0-2 years: 1; 3-5 years: 4; 6-10 years: 6; 11-15 years: 3; $\geq$16 years: 5
Gender Men: 14; Women: 5
Age 18-25: 3; 26-35: 10; 36-45: 4; 46-55: 2
Race White: 6; South Asian: 5; Asian: 2; Black or African American: 2; American Indian or Alaska Native: 1; Hispanic or Latino: 1; Jewish: 1; Multiracial: 1
Region US: 10; India: 2; Kenya: 2; Israel: 2; Germany: 2; China: 1
III-C Participants
After shortlisting repositories, we reached out to contributors who were employees of our company and who had contributed at least one commit to the repository in the previous 3 months. Additionally we aimed to recruit a diverse sample of participants by reaching out to participants belonging to different regions, genders, levels of seniority and with different levels of activity in the repository. 26 candidates declined or ignored our invitation to participate. We also recruited two participants based on recommendations from other invitees who were unavailable. In total, 19 participants accepted our invitation and completed the study. Participants demographics are presented in Table I. 18 of our participants reported having experience with Github Copilot Chat in their regular programming work, while 3 participants reported using Cursor.
III-D Protocol
Each study session was a meeting between the participant and the first author, conducted using a video conferencing tool. Sessions were scheduled for 60 minutes, with some variability based on configuration issues and participants’ availability beyond the scheduled time. At the start of each session, the study administrator asked the participant to fill out a pre-study questionnaire and consent form. Following this, the study administrator gave a tutorial of the Cursor IDE and its interface by working through a demonstration task to ensure that participants understood the tool, its usage, and its features. The study administrator instructed participants to think aloud and to focus on solving the issue, rather than assessing the AI, to create a more realistic environment. Participants were also instructed to try to use the AI for all their tasks, and make small manual changes if necessary – this instruction was added after some pilot study participants stopped using the agent altogether after switching to manual editing.
Participants worked on their tasks by remotely controlling the study administrator’s screen. We performed all the necessary set up for the participant’s repository on the study administrator’s system. We chose not to use the participants’ systems to eliminate the burden on participants on having to setup the Cursor IDE with their codebase. However, in one case, the participant worked on their own computer because there were configuration issues with the setup on the study administrator’s system and because the participant had previously installed and worked on the Cursor IDE.
Participants went through the list of open issues of the repository on GitHub to choose one to work on. [^2] Although we told participants that they did not need to prepare for the study, five participants planned to address a specific work item during the study. We allowed them to continue as they had planned. Once they completed an issue, the study administrator first asked the participant if they were satisfied enough with the fix to submit it as a pull request. If not, we asked them to keep working to get to that level of satisfaction (if possible within the scope of the session) before moving on to the next issue. We did not ask participants to create and submit a pull request since that would require them to enter their GitHub credentials onto the study administrator’s system. Nonetheless, so that participants could create a pull request with their credentials, we sent them the diff of their work as a patch after the session.
At the end of each session, participants filled out a post-study questionnaire containing questions about their perceptions of the tool.
III-E Data Collection
We collected several sources of data during the study, as described below.
TABLE II: Participant Activity Taxonomy
Activity Description
IDE Actions
Checkpoint Revert to an earlier checkpoint in the conversation
Accepted Code Change Accept code changes proposed by the agent
Rejected Code Change Reject code changes proposed by the agent
Stopped Agent Stop agent run prematurely
Communication
Prompt Crafting Writing a prompt in the chat window
Prompt Submitted Submitting a prompt in the chat window
Following Agent Live Waiting for/ following the steps of agent while it runs
Thinking/Verifying Explanations Actively thinking about and/or verifying textual information by the agent
Thinking/Verifying Actions Actively thinking about the actions taken by the agent to reach its final output (e.g. searching codebase for relevant files)
Thinking/Verifying Code Actively thinking about and/or verifying a proposed or accepted code change by the agent
Code Actions
Writing New Code Manually writing code that implements new functionality
Debugging/Testing Code Manually running or debugging code to check functionality (e.g. testing UI changes on local browser)
Editing Suggested Code Manually editing code changes suggested by the agent
Editing (Personally) Written Code Editing code written by a programmer that is not a suggestion by the agent for the purpose of fixing existing functionality
Non-Code Actions
Looking up issue Referring to issue description/ details
Looking up Documentation Checking an external source for the purpose of understanding code functionality or other information needed to help work on the issue (e.g. Stack Overflow)
Waiting for Terminal Waiting for terminal executions to complete
Deferring Thought For Later Programmer accepts suggestion without verifying/understanding it, but plans to verify it after
Thinking About New Code To Write Thinking about what code or functionality to implement and write (e.g. reading/navigating through the existing codebase)
TABLE III: Chat Trajectory Taxonomy
Code Description
Information Given to Agent in Prompt
Issue Knowledge Providing agent with instructions/knowledge about the issue based solely on original issue description, if any
Implementation Knowledge Providing agent with instructions/knowledge about the issue related to implementation/configuration/scope etc., not solely based on original issue description or environmental context
Styling Knowledge Providing agent with instructions/knowledge about the issue related to code styling/practices not solely based on original issue description or environmental context
Affirmation Affirming the follow up provided by the agent
Meta-details Informing the agent of plan-level heuristics or steps to follow in its response, not grounded in the codebase
Response Sought from Agent in Prompt
Change Explanation Seeking explanations of code changes made by the agent
Codebase Explanation Seeking explanations pertaining to the existing codebase
External Explanation Seeking explanations pertaining to external knowledge not rooted solely in the codebase
Refine Change Seeking corrected version of old code change
New Change Seeking new code changes
Testing Seeking to test existing/previously generated code
Agent Actions
New code change Add code changes to a file
Refine code change Refine previously made changes to a file
Run terminal command Run terminal command
Explain plan Textual description of multi-step plan to be performed by the agent
Explain codebase Explains details of existing code in the codebase
Explain old changes Provides textual explanations of code changes made in previous runs
Explain new changes Provides textual explanations of code changes made in current run
Search Web Searches the internet
Explain external Explains external details (usually based on web search)
Search codebase Searches through the codebase
Analyse files Reads/analyses files in the codebase
Follow Up Prompts the participant for follow-up queries (usually at end of response)
III-E1 Session recordings
To understand participants’ usage of the agent in the session, we used video, audio, and screen recordings to qualitatively coded both participant actions as well as the details of their interactions with the agent.
Participant Action Codebook
We coded participant actions by adapting the CUPS taxonomy, previously used for understanding common programmer actions when using AI code completion [21]. We modified this taxonomy to add states unique to collaborating with SWE agents, such as following the trajectory of the agent’s execution and backtracking to a previous point in the conversation (Figure 1a). Table II describes the specific codes we applied.
Chat Trajectory Codebook
We also coded the trajectories of the developer-agent interactions in our study by coding information on the context the participant provides to the agent (for example, files, code snippets, and external links) as well as the information they give to and seek from the agent in their prompts. To create our codebook, we performed open coding on two participants’ traces, then iterated on this codebook by annotating trajectories for two further participants to create the final version of our codebook, as described in Table III.
To validate each of our codebooks, one author coded a study session by applying codes to time segments across the session recording. These codes were then masked, and another author recoded the same session by applying codes to the corresponding time segments. We then computed Cohen’s $\kappa$, yielding inter-rater reliability scores of 0.92 and 0.89 for the participant action and chat trajectory codebook respectively, indicating near perfect agreement [44].
III-E2 Pre-Study Questionnaire
This questionnaire consisted of optional questions about participant demographics, as well as prior experience with AI tools. Prior experience questions were based on questions from previous studies of human-AI interaction [21] and developer-AI interaction [26]. We did this to explore how prior experience with AI tools might be related to participants’ use of the tool [26]. We based the list of the AI tools on those reported in the 2024 Stack Overflow Developer Survey [1]. We also included a question about participants’ perceptions of AI tools for programming in terms of usability [27], by adapting the Technology Acceptance Model questionnaire [45].
III-E3 Post-Study Questionnaire
This questionnaire contained questions about the usability of the tool; the perceived role of the agent in the development process [27]; the importance of the design features of the tool [26], where we drew features from the tool assessment in Section III-A; and an open-ended question for participants to describe any notable observations while using the tool.
III-F Analysis Approach
Based on relevance of actions to the immediate state of the chat with the agent, we interpret the presence and ordering of codes on participant actions at the prompt level (e.g., verifying code changes, stopping the agent) or at the issue level (e.g., writing code, running tests) and analyze codes on chat interactions at the prompt level. Since participants were instructed to think aloud, we do not analyze the durations of their activities, nor do we analyze the absolute counts of these activities to avoid biases toward participants who performed actions repetitively in shorter intervals. We also note specific utterances and scenarios that indicate communication barriers, reporting only those that recur across different participants.
While our results are largely qualitative, we include some numerical data throughout to give the reader a fuller sense of the trends that we observed. However, given the exploratory nature of the study and the relatively small sample sizes, we caution the reader against broad extrapolation or generalization. Congruently, we present findings descriptively, without applying inferential statistics.
Towards helping to ensure the validity of our findings, we sent an early version of our results to all participants as a member check [46]. Our study protocol has undergone privacy assessment and review by our organization. To aid in replication, qualitative codebooks and study materials can be accessed in our Supplementary Material.
Figure 3: Issue Timeline by Participant with Success Rates. Success Levels are detailed in Section IV-C.
III-G Limitations
One limitation of our study is that we recruited participants solely from one company. Participants’ interactions and perceptions of the tool are likely influenced by the corporate culture and organizational incentives at this company, which is actively promoting AI products and experiences. Likewise, we acknowledge our positionality as researchers in that company, and the inherent unconscious biases and incentives that this environment introduces in our interpretation of the study data.
Another limitation is small sample size (19 participants); we accepted this limitation so that we could perform time-consuming, thorough action and trajectory coding, which does not easily scale up to more participants.
Most participants in our study had not previously used Cursor or any other SWE agent, and thus our results mostly reflect early usage, rather than regular long-term use. Our study is also subject to observer effects; as participants knew that we were studying their interactions, they might may have used the tool differently than in a private setting. Likewise, our instruction that participants try to use AI for all their tasks means that, in the wild, developers would likely be more prone to abandonment and shifting to manual effort. Our choices of Cursor in early 2025 also has an effect on generalizability; since AI tools and models are constantly evolving, some findings reflect present-day limitations of agents and their underlying models.
IV Results
Figure 3 visualizes participants’ success and time spent as they submitted 269 prompts to the agent across 33 issues. Participants were satisfied with the final patch for 16 of these tasks, while 4 of these tasks were abandoned for external reasons (such as participants being unconvinced of the issue’s validity) for a total success rate of 55% $(16/(33-4))$. In this section, we first describe the patterns we observed when participants collaborated with the agent, followed by the communication challenges they faced in this collaboration. We then discuss how these patterns and other factors were associated with participants’ success.
IV-A RQ1: How developers collaborate with SWE agents
In this subsection, we describe the strategies participants applied when collaborating with the agent and the patterns this led to, ranging from high-level patterns related to delegation and iteration to lower-level patterns related to the type of information participants provided and the outputs they sought from the agent.
IV-A1 Delegation Strategies
We observed that participants broadly followed one of two approaches to delegate work –
In one approach (the one-shot strategy), participants provided the agent with the entire issue description and asked it to solve the issue. Ten out of 17 participants who worked on GitHub issues followed this strategy. This is a high risk, high reward approach:
  • If the agent is able to generate a comprehensive fix and verify its changes, the participant can resolve the issue with little additional effort.
  • If the agent is unable to generate a comprehensive fix or verify its changes, participants must then both manually understand the generated fix and then iterate on it to get to a valid fix. We observed that this iteration can be time-consuming, as the it can entail not only understanding the code change itself, but also understanding the reasoning behind the change, navigating between code changes, and understanding changes to test files.
In the other approach (the incremental resolution strategy), participants manually divide the task into sequential sub-tasks and ask the agent to solve each sub-task sequentially in separate prompts. When the participant is satisfied with the solution to one sub-task, they then ask the agent to work on the next sub-task. Although this approach is safer than the one-shot strategy because mistakes made by the agent can be reviewed and caught earlier, it requires more proactive involvement and communication from the participant to divide the issue into these sub-tasks. As evidence, participants who used this strategy used an average of 11.0 prompts per issue, compared to 7.0 for the one-shot strategy. Further, these participants were more likely to manually read existing code to enhance their own understanding while working on issues, doing so in 83% (15 out of 18) of the issues they worked on as compared to 60% (9 out of 15) for other participants.
One participant (P7), who had used Cursor before, explained that they preferred different strategies in different scenarios: “For an issue that is very contained, I trust it to come up with the changes itself, but for issues that are more involved with different parts of the codebase I feel have to point it in the right direction so that anything sensible comes out at the end”. However, we also observed that these approaches are not mutually exclusive; when delegating the entire issue to the agent via the one-shot strategy, an incomplete solution may require the participant to prompt the agent with extra steps via incremental resolution. P3 noted how they dynamically changed their level of involvement during the collaboration after starting with the one-shot strategy: “The AI agent was able to generate good structure… once more context was needed, it committed errors that need more handholding and additional time to be spent on and guide it”.
Figure 4: Post-study questionnaire responses on Agent Perception
IV-A2 Knowledge Transfer
When working with the agent to solve issues, we observed that participants provided two different types of information to the agent to help it make the desired code changes:
  • Contextual knowledge from the issue description and the environmental context (such as test logs). Eg: “There are build errors in your solution”
  • Expert knowledge relating to code implementation or convention that cannot be observed directly from context or the issue description and is based on the developer’s prior knowledge of the repository. Eg: “Usually we don’t set an upperbound. I think 3.10 should be fine now”
Participants’ patterns of providing such knowledge to the agent mirror those of software engineering managers. Kalliamvakou and colleagues [47] report that when assigning work, great managers in software engineering leave the implementation details to the engineer, providing actionable feedback when required. Analogously, participants tended to provide less expert knowledge to the agent up-front, and more as feedback; participants provided expert knowledge in 49% of prompts that asked the agent to generate a new code change, compared to in 65% of prompts that ask the agent to refine a code change (i.e., iterate upon a previously generated change). This relationship was even stronger for the 12 participants in managerial roles – they provided expert knowledge in 45% of their prompts seeking new changes versus 75% seeking to refine changes. As shown in Figure 4, the analogy is reflected in the post-study questionnaire, in which participants report viewing the agent more as an assistant (at a lower level in the organizational hierarchy) than as a colleague (at the same level) or advisor (at a higher level).
While participants usually tended to provided expert knowledge in feedback prompts rather than up-front, we found that experience factors moderated this relationship. Participants who had more familiarity with the repository tended to provide expert knowledge more often – participants with over 10 commits to the repository in the four months leading up to the study provided such insights 70% of time when seeking any code changes as compared to 48% for other participants. Similarly, participants who had previous experience with Cursor also provided more expert knowledge, doing so in 81% of prompts seeking code changes. These participants also tended to provide expert knowledge more often as feedback – they did so for 89% of prompts trying to refine a code change as compared to for 67% of prompts when asking for a new code change. Further, participants applying the one-shot strategy provided expert knowledge less often (in 47% of their prompts seeking code changes) as compared to those applying the incremental resolution strategy (did so in 64% of their prompts seeking code changes).
IV-A3 Tasks Requested
We observed that participants not only requested the SWE agent to make code changes to resolve the the entire task specified in the GitHub issue but also requested assistance in sub-tasks corresponding to other activities of the software development process. While 50% of participants’ prompts requested code changes, 27% of the prompts sought explanations of details related to the existing codebase (code comprehension), 16% of these prompts sought to run tests and 11% of these prompts sought explanations related to changes made by the agent (code review). Note that the above categories are not mutually exclusive. This distribution of prompts reflects participants’ perceptions of the agent as reported in the post-study questionnaire (Figure 4). Participants viewed the agent the most as a content generator; prompts requesting code generation were most common. Since ‘reference guide’ was a close second, participants also often used prompts to ask the agent about the codebase. On the other hand, participants did not view the agent as a reviewer, and were thus less likely to ask it to run tests or review its changes.
IV-A4 Manual Actions
Participants’ perceptions of the agent shaped not only the types of prompts they provided the agent, but also the role the participants themselves took in the collaboration, tending to work on verification (debugging and testing) more than on content generation (writing code). Since participants reported not thinking of the agent as a code reviewer (Figure 4), they often took the responsibility to manually debug and test code upon themselves, doing so in 21 out of 33 issues by running tests or examining execution logs. In the case of bugs concerning UI changes, it was also easier for the participant to test changes manually by simply observing the current state of the UI rather than uploading images to the chat for the agent to interpret.
On the other hand, since the participants tended to view the agent more as a content generator, they were more likely to let the agent handle all the code changes to be made, manually writing code in only 14 issues. Further, in 10 of these issues, participants’ code interventions were limited to editing code suggested by the agent rather than adding new functionality.
Some participants expressed desire for AI tools to provide more support in reviewing and verification actions such as localizing, debugging, and testing, since these aspects of the software development process were often critical to resolving the issue. P12 expressed that that the agent was not particularly helpful in localizing the issue, mentioning that “I think the main part [of solving the issue] was finding where to put the code, so it might have been marginally faster to do it myself”. Similarly, participants mentioned the challenges they faced when collaborating with the agent for debugging; for instance, P8 responded in the post-study questionnaire that the agent was “not too great or useful at the debugging and converting-a-repro-into-a-test loop”, and P11 responded that “It’s puzzling to me why [the agent is] so good at generating code but bad at dealing with basic environmental commands”. Participants’ manual actions thus reflected not only their perceptions of the agent, but also a forced response to the constrained capabilities of the agent as a collaborator.
IV-A5 Reviewing Outputs
The agent provided output to communicate with participants in three ways:
  • Execution Trajectory: a trace of the agent’s work that appears while the agent is working, including context gathering steps such as codebase searches and files read (Figure  1b), along with actions taken such as terminal commands executed and files edited. (the bulk of Figure 1).
  • Change Explanation: at the end of a trace, typically the agent summarized the change it made to the participant’s codebase (Figure 1c).
  • Change Diff: inside the participant’s editor, the agent provides a number of inline diffs as proposed changes that the participant can accept or reject (Figure 2).
We observed that participants tended to first track the agent’s execution live as-it-happened before reviewing explanations and diffs, with a preference for reading diffs. While Cursor enables its users to do manual tasks in parallel to its own execution, participants instead followed the agent’s execution as it happened in real time for 84% of their prompts. In some cases (23% of prompts), this was sufficient validation for participants, as they did not further verify the diffs or agent explanations. Usually, this live analysis was at least sufficient for participants to understand the agent’s context gathering steps; after following the agent’s execution live, they only reviewed this context gathering process for 7% of the agent’s responses. However, in the majority of cases (specifically, after 75% of the agent’s responses), participants reviewed diffs or explanations after following its execution in real time.
When reviewing code outputs, participants preferred directly verifying the generated code rather than reviewing code explanations. While participants followed the agent’s execution in real time for 95% of the responses that included code changes, these code changes often required further verification – participants reviewed diffs after 67% of such agent responses, while they only reviewed explanations after 31% of these responses. Of these 31%, participants verified both the diff and the explanations after 23% of such responses. This preference to review code is also reflected in the post-study questionnaire, in which participants ranked the ability to see detailed explanations of code changes as the least important feature among the different features of the agent (Figure 5). Such manual verification was usually sufficient for participants to have clarity on the agent’s code changes – they only asked for explanations of generated changes in 16% of their follow-up prompts.
Participants who had previous experience with Cursor tended to focus even less on explanations than their peers, without any apparent impact on their success (Section IV-C2). Like other participants, they followed the agent during 95% of responses containing a code change, and further verified diffs in 66% of these responses. However, they only read explanations after 18% of agent responses with code changes, as compared to 35% for other participants.
Figure 5: Post-study questionnaire responses on Agent Features
IV-A6 Handling Incorrect Outputs
To address incorrect code generated by the agent, participants in our study preferred to refine the agent’s changes by iterating on these changes rather than abandoning them and starting over. Such refinement was required more often when working on feature requests – participants used an average of 11.2 prompts when working on feature requests as compared to 6.3 when working on bugs. Overall, participants used an average of 8.2 prompts per issue, out of which 52% of prompts that sought code changes sought refinements to previous changes by the agent. Thus, the agent’s first outputs were often unsatisfactory and required further modifications. Despite this, participants only rejected code changes 10% of the time after the agent generated code changes. Further, while the agent allowed for participants to reset the chat to earlier points in the conversation (Figure 1a) – a feature we mentioned in the warm-up task – participants only used this feature in 15% of the issues they worked on. Additionally, participants only terminated the agent’s execution before completion (Figure 1e) in 11% of prompts. Thus, participants did not tend to backtrack by stopping or resetting the agent, instead preferring to iterate on imperfect changes.
Despite participants’ proclivity to iterate on incorrect outputs, some participants indicated that they did not think that iterating was always a good strategy. P14 mentioned of the agent in their post-study questionnaire that “if you continue to ask it to fix issues that it hallucinated on you might end up wasting time following in on the wrong path. It would be better to just explain the task in more detail or revert back to a previous breakpoint.” Similarly, P6 noted that the agent “was overly generative, not understanding the implications of previously written bad code, and doubling down by adding more code”.
IV-B RQ2: Developer-Agent Communication Barriers
IV-B1 Lack of tacit knowledge
The agent was sometimes an ineffective collaborator because it lacked tacit knowledge, a type of knowledge that is gained through personal experience, rather than through explicit documentation. Rus and Lindvall [5] show that tacit knowledge comprises the majority of software engineering knowledge. The lack of such knowledge can make it difficult for new collaborators (such as the agent) to work effectively as they lack the knowledge that developers who are experienced with the repository, such as the participant, may have. Further, unlike expert insights into implementation details, it is difficult for the agent to learn tacit knowledge, as it is difficult to know when and how to communicate such information.
For example, in the case of P7 (who had prior experience with Cursor), the participant was able to interpret a temporal cue but unable to communicate their finding to the agent. When a terminal command running unit tests that was started by the agent seemed to be taking a long time to finish execution, the participant realized based on their experience with the repository that the tests were failing. However, rather than stopping the agent’s execution and explaining this in a prompt, they chose to work on the issue manually while the agent came to this realization by itself. In other cases, the agent suffered from a lack of social context. For example, the agent once offered an explanation of the codebase that seemed to contradict comments on the issue by a maintainer of the repository, making participant P8 very skeptical of the agent’s correctness. Prior knowledge of this context might have led the agent to question its assumptions before generating such an explanation.
IV-B2 Unsolicited Actions
We observed that the agent was sometimes excessively proactive in making changes to solve the issue, making changes beyond the prompt’s scope. We found that the agent made such unsolicited changes after 38% of participants’ prompts that did not seek code changes. In some cases, this was productive. For example, P5 mentioned “[the agent] anticipated what I wanted to do”. In other cases, this led to miscommunication. For example, when the agent made changes beyond what the prompt specified, P16 was forced to reject the change and write a follow-up prompt to specifically instruct the agent to only make the requested change. Another participant mentioned that the agent “always tries to generate new code, even though the right thing to do here might be [something else]”.
We further observed that the agent ran terminal commands in 10% of prompts that did not request to run any. As further evidence that these commands were unwanted, participants were nearly three times more likely to stop the agent prematurely when its response included terminal commands (61% vs. 21%). Unlike in code changes, where the agent does not apply the changes until proposed and can undo changes, terminal executions can permanently change the state of the environment and the codebase. For example, while P12 was reviewing a test log file, the agent ran some new tests on the terminal, even though this was not instructed in the prompt. This led to the test log being deleted, derailing the participant’s work. Similarly, P11 had to abandon an issue due to the agent corrupting the code environment by running a long series of terminal commands.
IV-B3 Synchronicity
Parallel actions by the agent and the user sometimes led to challenges stemming from the participant’s unawareness of the actions being performed by the agent. While in theory simultaneous collaborations on the same codebase allowed for participants to never be blocked from directly working on the code, this did not always work out in practice. When P9 tried to work on code alongside the agent, it led to the agent duplicating a part of the fix, forcing the participant to reject the changes. To avoid such situations, some participants preemptively prevented themselves from working on the codebase. P10 expressed that “I want to make sure that the code changes are clear and I can see the diffs clearly so I’ll wait for it to stop thinking”. While the agent was iterating on a fix and seemed to be going down the wrong path, P11 noted similar concerns: “It is doing something which is no longer helpful… I am scared it will break something, and so I will wait for it to finish what its doing”. Pu and colleagues [34] similarly reported that proactive AI assistance can make users uncertain about what actions they can perform without disrupting the AI, and suggest addressing this by limiting the AI’s actions to those that do not interfere with the user’s immediate workflow.
IV-B4 Verbosity
Participants often sought and read explanations by the agent, doing so after 39% of the agent’s responses. However, several participants (P6, P7, P13, P15, P18) expressed frustration with the verbosity of such explanations. While P15 noted that “Most of the time when I’m coding, I don’t need detailed responses, only to see the changes made”, P16 mentioned of a response that “It’s giving me a lot of gibberish… I don’t need all this analysis”. Sometimes, verbosity even led to miscommunication. In the case of P15, due to the verbosity of the agent’s response, they did not notice that the response contained a documentation link that they were looking for, and instead searched the web manually for the link. Thus, while participants engaged with the agent’s explanations, these explanations did not always provide participants with the level of details they needed for effective communication. This observation echoes those made by Bansal and colleagues [32]
IV-B5 Sycophancy
Sometimes, the tendency of the agent to agree with whatever the participant mentioned in their latest prompt led to a mistrust of the agent. In particular, when the agent produced a fix to an issue and an explanation for that fix, participants were hesitant to trust the agent when in subsequent responses it offered a seemingly contradictory explanation for the same issue. As P15 mentioned, “If I say that ’your change is wrong’, it will revert that change. That makes me lose confidence in it”. Similarly, P16 expressed in the post-study questionnaire that the agent “can be confused easily and mislead the developer so it can’t be trusted blindly”. Bansal and colleagues [32] similarly note that consistency in the agent’s outputs in critical to establishing trust in the agent. P18 applied an alternate approach that avoided such issues, explaining that “I tend to treat it like a human, where I don’t want to give it the answer, I want it to think about it, and come to a conclusion itself. It might have insights I don’t know myself”. They did this by asking the agent suggestive questions (e.g., are there any negative consequences of…) rather than instructing the agent. This allowed the agent to retrospect while making refinements rather than blindly changing its assumptions based on the participant’s latest instructions. This approach worked well for the participant - they were able to solve the issue as well as find additional related problems with the code (unnecessary type checking) that they did not know about previously.
IV-B6 Overconfidence
Another aspect of the agent’s response that caused participants to mistrust it was the agent’s conviction that the changes it made in its latest response correctly solved the issue. As P9 noted in the post-study questionnaire, “It never stopped thinking it knew how to solve the problem even though it was wrong”. Even when the agent generated correct changes and explanations, its confidence in the alignments of these outputs to the issue sometimes led the participant to be hesitant of the agent. P18 noted about a generated fix that “it works, but its not really what I want”. Similarly, upon reviewing the agent’s description of files relevant to addressing the issue, P10 mentioned, “These things are true, but not necessarily things I want to modify. Khati and colleagues [48] similarly report that AI trust metrics for developers include not only correctness, but also alignment with their intentions.
Participants were particularly hesitant when the agent made many changes. Participants stopped the agent’s execution prematurely in 39% of the responses in which the agent performed more than 3 actions (such as terminal commands or code changes), as compared to only 9% when the agent performed 3 or fewer actions. Participants sometimes expressed their concerns over losing control of the agent – while P10 mentioned that “I don’t feel super comfortable when it makes edits and I can’t track exactly what and where”, P14 noted about an issue that they were not able to successfully complete that “the problem I had with the agent is that it took over, I had to force delete things and it created a mess”.
IV-B7 Ineffective Follow-Up Suggestions
The agent offered textual suggestions for follow-up prompts in 62% of its responses. These suggestions were usually not effective in advancing the conversation– participants replied to only 7% of such follow ups in their next prompt. One reason for this may be that these suggestions were often non-specific (e.g.: Figure 1d). When participants did reply, the suggestions were more specific (e.g.: Would you like me to show you how to remove the manual tracking code and let RichEdit handle double-clicks naively?). Further, while participants often needed refinements to code changes, the agent’s follow ups instead would often suggest moving forward to other steps in the issue resolution process. Thus, participants may also not have responded to follow ups because they did not always reflect the true state of the solution.
IV-C RQ3: Factors Associated with Success
We considered participants to be successful in solving the issue if they were completely satisfied with the fix so as to submit it as a pull request (Complete Success, 10 issues) or if they were satisfied with the generated patch but needed to run some configuration steps or tests outside the scope of the study session before submitting the change as a pull request (Patch Success, 6 issues). Participants made progress but were unable to fully complete 10 issues (Incomplete), did not make any progress in 2 issues (No Progress), and abandoned 1 issue due to an inability to make any progress towards the fix (Abandon Fix). Although the total number of issues that participants worked on (N=33) is too low to conclusively reason about participants’ success in resolving issues, we observed that two main factors were associated with success.
IV-C1 Collaboration Strategies
Participants were more likely to be successful in collaborations with the agent when they actively and regularly contributed to the problem-solving process.
For example, participants were more likely to successfully resolve issues when they communicated more with the agent – the average number of prompts used for successful issues was 10.3 as compared to 7.1 for unsuccessful issues. Refining the agent’s outputs was also a factor in success. Only in 19% of the successful issues did participants not request any code refinements from the agent. In addition, participants were only successful in 30% of the issues in which they did not ask for code refinements.
Notably, participants who applied the one-shot strategy (Section IV-A) were less successful in solving issues – they only succeeded in 38% of the issues for which they provided the agent with the entire issue description. On the other hand, participants who applied the incremental resolution strategy were successful in 83% of the issues they worked on. Similarly, participants who provided the agent with expert knowledge (Section IV-A2) tended to perform better than those who simply provided the agent with environmental context and thus left the entire task implementation to the agent without providing actionable feedback. Participants succeeded in 14 out of 22 (64%) issues in which they supplied such expert insights to the agent, as compared to only 2 out of the 7 (29%) issues in which they did not provide any expert knowledge.
Beyond providing insights to the agent, we also found that participants were more successful when they contributed to the code directly rather than just relying on the agent to generate the fix. Participants were successful in 79% (11 out of 14) of the issues in which they wrote some code manually, as compared to in 33% (5 out of 15) of the issues for which they did not write any code manually. This result did not extend to manual debugging and testing – participants who manually worked on debugging or testing code for an issue were successful in 53% (10 out of 19) of such issues, as compared to 60% (6 out of 10) for issues in which they did not perform manual debugging or testing.
IV-C2 External Factors
Participants’ success was influenced by external factors relating to the tool and the task, such as prior experience and nature of the issue. The 3 participants with previous experience with Cursor had more success (3 out of 4 issues resolved), compared to others’ 52% (13 out of 25) success-rate.
Participants were most successful solving bugs that only concerned user interface changes (5 out of 5 issues resolved), followed by refactoring/variable renaming issues (2 out of 3 issues resolved), then feature requests (4 out of 8 issues resolved), and finally bugs (5 out of 13 issues resolved). These results suggest that participants may have struggled to collaborate with the agent on non-code generation tasks such as localization and debugging that are critical to bug resolution.
While benchmarks suggest that autonomous agents perform differently when working on code in different programming languages [49], our human-in-the-loop study suggests that this was not an important factor in participants’ success. Participants had similar success rates – succeeding about half the time – across different languages (Python: 4 of 9 resolved successfully; TypeScript: 5 of 8; C++ 3 of 6; Cmake: 1 of 2; and Java 0 of 1), with the exception of C# (3 of 3).
V Discussion
V-A Collaborating Actively across the Development Process
A recent blog post by Anthropic [50] discusses the prevalence of the ‘feedback loop’ interaction pattern, in which users collaborate with agents to work on code tasks simply by providing the agent with feedback from the environment. In our study, we observed a similar interaction pattern, as many of participants’ prompts requesting code changes consisted only of environmental context or details of the original issue description. However, our results suggest that such interactions are less effective than those in which developers provide their own expert knowledge to the agent (Section IV-C1). Thus, when collaborating with agents, developers should not restrict their role to ‘human routers’ [51] that simply provide environmental feedback, but rather actively contribute to the problem-solving process to achieve better results. Since participants often took a managerial role in these collaborations (Section IV-A2), future work could explore how agents can adopt the communication strategies and behaviors exhibited by high-performing subordinate engineers.
Further, our results also suggest that developers should not restrict their active involvement to the parts of the software engineering process in which the agent is not particularly helpful, such as debugging and testing. Even though participants primarily used the agent to generate code changes, they were more successful when they collaborated with the agent by also manually writing code. On the other hand, while participants expressed misgivings about the agent’s ability to help with debugging (Section IV-A4), manually debugging and testing code did not appear to improve their likelihood of success.
Thus, developer-agent collaborations were most successful when both the developer and the agent actively contributed to issue resolution. In line with prior findings [33], as the agent was not as supportive in debugging and localizing, it was often unable to actively contribute to this part of the issue resolution process, and thus participants struggled to successfully resolve bugs. While software engineering benchmarks have clear, well set up tasks and environments for agents to work on, real development can involve messy environmental setup and debugging steps before the developer is ready to make code changes. However, agents often fail at these tasks, as reflected in their worse performance on more realistic benchmarks [4]. For SWE agents to be useful for developers in their day-to-day tasks, they must thus empower users across all the steps of the software engineering process, especially the messy and ambiguous ones.
V-B Calibrating Output Scope
AI agents are designed to solve complex issues autonomously, and have the tools to do so [2]. However, we found that this design can backfire – in our study, participants who applied the one-shot delegation strategy were often unsuccessful (Section IV-C1). When the agent tries and fails to successfully implement complex changes, it is difficult for the developer to debug where the problems in the buggy patch arise from. Further, real-world complex issues often require expert knowledge that only an experienced developer might have. Yet, knowing what information could be relevant may be difficult for developers to foresee, and providing all such necessary information in a single prompt might be tedious for developers, particularly in the case of complex issues. Thus, interactive agents should leverage the presence of the user as a collaborator by internally calibrating the scope of their outputs when working on an issue. This could enable the developer to provide expert insights when needed, and effectively investigate any problems with the agent’s output. Further, such calibration enables the agent to have a clear idea of potential next steps and be aligned with the state of the solution, engendering actionable and relevant follow-up suggestions (Section IV-B7). Bajpai and colleagues [25] similarly found that in-IDE AI copilots are more effective when debugging if they focus on gathering information and communicating with the user before suggesting a fix. Such a pipeline, as also supported by [52], has the advantages of the incremental resolution strategy without requiring additional effort from the user to manually analyze the issue or intervene in the fix unless necessary.
V-C Collaborating through Discussion, Not Sycophancy
Friction arising from the agent’s lack of tacit knowledge is difficult to anticipate for both the agent and the developer (Section IV-B1). This friction may occur regardless of how much context the agent gathers, which suggests that agents may always produce incorrect patches in some cases.
In contrast to pair programming, where developers spend considerable time discussing code logic and implementation details [53], the agent rarely engaged in problem-solving discussions. Instead, despite initial confidence in its changes, the agent immediately agreed with whatever the developer said in the latest prompt. While participants using incremental strategies were more successful, the agent contributes little to the solution quality in this paradigm since it simply follows participant instructions.
However, when participant P18 invited the agent to challenge their opinions, they uncovered novel insights about the existing codebase, a synergy highlighted in recent surveys of LLM-backed SE agents [2]. This suggests that agents designed to challenge rather than merely obey the developer – by engaging in substantive discussion – might not only increase trust and improve solution quality, but further promote developer growth through critical thinking [54].
VI Conclusion
In this paper, we’ve seen that developers can collaborate with a SWE agent to solve real-world software engineering tasks, but also that such a collaboration is fraught and that success is not guaranteed. While the capabilities of large language models and of the agents that build upon them expands at a dizzying rate, the capabilities of the humans who work with them are relatively static. The strategies, challenges, and communication patterns we document here are therefore likely to remain relevant even as SWE agents rapidly evolve. By understanding and addressing these enduring human factors, we hope this work contributes to a future where developer-agent collaboration is both effective and empowering.
A. Pre-Study Questionnaire
We asked the participants to answer the questions below before the study:
  1. 1.
    With which gender do you most identify?
    • Woman
    • Man
    • Non-binary / gender diverse
    • Prefer not to say
    • Other
  2. 2.
    What is your age?
    • 18-25
    • 26-35
    • 36-45
    • 46-55
    • 56-65
    • 66+
    • Prefer not to say
  3. 3.
    With which race/ethnicity do you most identify? (Select all that apply)
    • American Indian or Alaska Native
    • Asian
    • Central Asian
    • East Asian
    • South Asian
    • Southeast Asian
    • Black or African American
    • Native Hawaiian or Pacific Islander
    • White
    • Hispanic or Latino
    • Multiracial
    • Prefer not to say
    • Other
  4. 4.
    Which best describes your programming experience?
    Scale: Multiple Choice
    • 0 to 2 years professional programming experience
    • 3 to 5 years professional programming experience
    • 6 to 10 years professional programming experience
    • 11 to 15 years professional programming experience
    • More than 16 years professional programming experience
  5. 5.
    When coding, how frequently do you use the following tools to chat with AI?
    Scale: Never, Occasionally, Regularly
    • ChatGPT
    • GitHub Copilot Chat
    • Codeium Windsurf
    • Cursor
    • Tabnine
    • Claude
    (Scale for each: Never / Occasionally / Regularly)
  6. 6.
    If you have previously used any tools other than the ones mentioned above to chat with AI, please specify them below.
    Open-ended response
  7. 7.
    When coding, how frequently do you use the following tools’ AI-generated code completions?
    Scale: Never, Occasionally, Regularly
    • GitHub Copilot
    • Windsurf Cascade
    • Cursor AI
    • Tabnine
    • Blackbox AI
  8. 8.
    If you have previously used any AI tools other than the ones mentioned above for AI-generated code completions, please specify them below.
    Open-ended response
  9. 9.
    How frequently do you use AI developer tools?
    Scale: Never, Occasionally, Regularly
    • to write new code
    • to write documentation for existing code
    • to modify existing code
    • to explain unfamiliar code
    • to debug code
  10. 10.
    Please rate your agreement with the following statements. AI tools for fixing bugs:
    Scale: Not at all, Moderately, Extremely, Don’t Know
    • are easy to use
    • can understand your requests
    • can provide high quality responses
    • can help you to write better code
    • can help you to write code more quickly
    • are enjoyable to use
B. Post-Study Questionnaire
We asked the participants to answer the questions after the session with Cursor:
  1. 1.
    Please rate your agreement with the following statements. AI tools for fixing bugs:
    Scale: Not at all, Moderately, Extremely
    • are easy to use
    • can understand your requests
    • can provide high quality responses
    • can help you to write better code
    • can help you to write code more quickly
    • are enjoyable to use
  2. 2.
    To what extent did you view the AI tool as:
    Scale: Not at all, A little, Somewhat, A great deal
    • A reference guide
    • A content generator
    • A problem solver
    • A collaborator
    • A colleague
    • A coach
    • An advisor
    • An assistant
    • A reviewer
  3. 3.
    How important were these aspects of working with the AI tool?
    Scale: Not Important, Moderately Important, Very Important
    • Detailed textual explanations of code changes
    • Ability to see the steps that led the AI to the relevant files for the query
    • Ability to follow the steps the AI took as they happened in real time
    • Ability to reject and/or accept code changes proposed by the AI
    • Refining the outputs given by the AI by continuing the conversation
    • Ability to undo actions by resetting to the state of earlier queries in the chat
    • Ability of the AI to use the currently open file as context for its response
    • Ability to refer to files or code blocks for the AI to use as context
  4. 4.
    Did anything stand out to you about the experience of using the AI tool? For example, was anything good, bad, surprising, or notable?
    Open-ended response
C. Study Protocol
Following is the study protocol containing details of the tutorial task and guidelines for the tasks participants performed that the study administrator followed across all sessions:
Introduction
Hi, good morning/afternoon/evening. Thanks for taking time out to participate in this user study. Before we begin, could you fill a short pre-study survey? I’ll send over the link to the form, and you can fill it on your system. Your participant number is ____
Participant fills the pre-study questionnaire
Today, you will be using the AI in Cursor IDE by remotely controlling my system to work on issues in the ____ repository. You can choose to work on any issues that you feel you could solve by making direct edits to the codebase. The goal is for you to create a solution patch for the issue that you would feel confident attaching your name to and submitting as a pull request for review. If you get done solving the issue, we’ll choose another issue for you to work on. The session will be around an hour long, and you will fill out a short post-session survey at the end. As mentioned in the survey, I’ll be recording the meeting. Do you have any questions?
Great, let me share my screen.
Warm-up Task
Let’s go through a quick demo task so you are familiar with how the IDE and its AI features work. Cursor is a fork of VSCode, so it uses most of the shortcuts and functionalities present in VSCode. Today, we’ll be using the AI tool Cursor Agent. You can chat with Cursor Agent and ask it to implement code changes for you. Cursor Agent can read and analyse all any files in the IDE, make code edits, and also run terminal commands. Let’s try this out. I have cloned a GitHub repo (link) for a CLI-based python game that involves sorting balls into different containers. Let’s try out this game on the terminal by running python3 cli-ball-sort.py. Great, so as you can see, there are 5 different containers and 3 balls to sort. So now, let’s say we want to use Cursor Agent to modify the game. On the chat window on the right, I’ve entered a prompt (Prompt entered: change the number of containers to 6, the maximum number of balls in each container to 5, and test out the changes) that asks to change the number of containers to 6 and the number of balls to 5. Cursor uses the open file as context for its response, though it can also search through the codebase for other files as we’ll see. You can also manually add context to give to Cursor. For example, I can add some lines of this file (customizer.py) to the context. I can also remove context (remove those lines). Let’s try giving it our prompt.
Cursor Agent runs and makes edits
As you can see, the AI searched through the different files and made edits. You can now approve or reject these edits. Cursor has also run the game through the terminal, and it seems to be working. Let’s check how it did by running it ourselves.
Run the game on terminal
Great, since these changes look right, let’s approve them. Aside from rejecting these changes, you can also stop the agent at any point in its execution and restore to an earlier point in the conversation (demonstrate where this button is). Do you have any questions about Cursor?
Great, let’s get started with the study. After you’re doing solving an issue, I will save the git diff of the changes you made and revert the repository back to the original state. I’ll email you these changes to you after the session.
Main Task
The repository has been set up in Cursor along with the all the required installations on a new conda environment. As in the demo, you can use the terminal.
You are also free to access the internet. I have a browser window open with the open issues in this repository (open the browser window), you can use this window to access any resources you need.
Please try to think aloud as you go though your process. Try to use the AI for all your tasks in the study, but feel free to make small manual changes if necessary.
Just to reiterate, your objective in this session is to solve the issue efficiently to the best of your ability, not to evaluate or observe how well the AI can solve the issue.
Let me give you remote control of my screen. You can now look for an issue to start solving with Cursor Agent.
Reminders
Before moving on to a new issue, confirm if the participant is 100% happy with the change, and they can attach their name on the patch when submitting a PR. If not, ask them to keep working on the issue till they get to this point. Only move on in this case if it is not possible for them to reach this level of satisfaction with the issue in the scope of the study.
Choice of issue:
1st issue: Entirely participant’s choice
Subsequent issues: Work with participants to choose issues varying in difficulty and domain (e.g. feature vs bug) depending on how successful the first issue was.
Prompts to remind the participant to think aloud:
  • Can you explain what you’re trying to do right now?
  • Could you talk about what your thought process right now?
  • What are you thinking about/ trying to do?
  • Tell me a little what you’re doing
  • Could you explain how what you’re doing is relevant to the task?
Post-Task
Thanks for participating! The session was very insightful. I’ll send over the link to the post-study survey so you can fill it out on your system.
Participant fills the post-study questionnaire
Great! Before we end, do you have any questions?
Okay, then once again, thanks for participating! Have a nice day!